# ToolTrain: Tool-integrated Reinforcement Learning for Repo Deep Search

### v1

@@ Abstract
Issue localization, the process of identifying code locations that need modification to resolve software issues, is a critical yet challenging task in software development. The semantic gap between natural language issue descriptions and faulty code requires complex multi-hop reasoning through code dependencies. Existing LLM-based agents attempt to address this by integrating repository retrieval tools, which poses a higher demand for LLMs to effectively utilize various tools during multi-step reasoning for issue localization. To tackle this challenge, we present ToolTrain, a two-stage tool-integrated training framework combining rejection-sampled supervised fine-tuning and tool-integrated reinforcement learning to enhance LLMs’ ability to use retrieval tools for issue localization. Experimental results show that ToolTrain-trained models achieve state-of-the-art performance, with our 32B model even surpassing Claude-3.7 on function-level localization. The results also show that improved localization performance translates to better end-to-end issue resolution performance. This further demonstrates that training for issue localization is a viable and effective strategy for improving automated software development.
@@ 1 Introduction
Despite this promise, LLM-based issue localization faces challenges. First, issue descriptions and their corresponding faulty code often lack direct semantic connections. Like human developers, LLMs must perform multi-hop reasoning through code dependencies, tracing function calls and variable usages across multiple files to connect high-level issue descriptions with low-level fault locations. Second, enabling LLMs to directly locate faulty code across the entire repository for such multi-hop localization remains difficult, as it is impractical to input entire repositories into LLMs due to context limitations (Radford et al., 2019).
Recent research has addressed this challenge by integrating repository retrieval tools into LLM-based agents (Chen et al., 2025; Yang et al., 2024), allowing LLMs to use various tools (e.g., searching specific function by keyword, listing functions that call a certain function) to dynamically explore repositories. However, this approach places higher demands on LLMs: beyond multi-step reasoning, they require sophisticated tool-calling abilities to explore code effectively rather than walking aimlessly through the repository. Current LLMs often struggle with these high-demand requirements, making incorrect tool calls or failing to maintain coherent reasoning chains throughout the exploration process.
@@ 2.2.1 Rejection-Sampled Supervised Fine-tuning
2.2.2 Rule-based Reinforcement Learning
2.3 Motivation
3 Approach
3.1 Issue Localization Agent
@@ 4.2 Metrics
To systematically evaluate the effectiveness of RepoSearcher with ToolTrain-model, we conducted a multi-dimensional evaluation using multiple metrics.
Recall@k
MAP
[…]
MRR
[…]
nDCG@$k$
[…]
%Resolved
@@ 7 Conclusion
In this paper, we improve the issue localization performance of LLMs by introducing ToolTrain, a tool-integrated training approach to enhance their tool-use and reasoning. Our framework first utilizes supervised fine-tuning to warm up the model with our lightweight agent, RepoSearcher, and then employs tool-integrated reinforcement learning to teach the model how to effectively navigate code repositories using various tools. We apply ToolTrain to open-source models, and evaluate them on issue localization tasks. The experiment results show that ToolTrain trained model achieve state-of-the-art performance among same-size models, and even surpassing leading models like Claude-3.7 on specific tasks.

### v2

# Tool-integrated Reinforcement Learning for Repo Deep Search

@@ Abstract
Issue localization, the process of identifying code locations that need modification to resolve software issues, is a critical yet challenging task in software development. The semantic gap between natural language issue descriptions and faulty code requires complex multi-hop reasoning through code dependencies. Existing LLM-based agents attempt to address this by integrating repository retrieval tools. However, this transforms issue localization into a demanding task we call Repo Deep Search, which requires the LLM to effectively utilize various repository retrieval tools throughout a multi-step reasoning and navigation process. To tackle this challenge, we present ToolTrain, a two-stage tool-integrated training framework combining rejection-sampled supervised fine-tuning and tool-integrated reinforcement learning to enhance LLMs’ ability to use retrieval tools for issue localization. Experimental results show that ToolTrain-trained models achieve state-of-the-art performance, with our 32B model even surpassing Claude-3.7 on function-level localization. The results also show that improved localization performance translates to better end-to-end issue resolution performance. This further demonstrates that training for issue localization is a viable and effective strategy for improving automated software development.
We shape our tools, and thereafter our tools shape us. —Marshall McLuhan
@@ 1 Introduction
LLM-based agents (Chen et al., 2025; Yang et al., 2024) allow LLMs to use various tools (e.g., searching specific function by keyword, listing functions that call a certain function) to dynamically explore repositories. However, this approach places higher demands on LLMs: beyond multi-step reasoning, they require sophisticated tool-calling abilities to explore code effectively rather than walking aimlessly through the repository. We refer to this complex, sequential navigation and discovery task as Repo Deep Search. Current LLMs often struggle with these high-demand requirements, making incorrect tool calls or failing to maintain coherent reasoning chains throughout the exploration process. Inspired by recent advancements in information retrieval where Reinforcement Learning (RL) agents are trained for multi-step Deep Search (dee, 2025; Jin et al., 2025), we investigate whether this paradigm can be extended to Repo Deep Search. Specifically, we explore using RL to enhance an LLM’s autonomous navigation within a code repository, aiming to significantly improve its performance on what we define as Repo Deep Search for issue localization.
@@ 2.2.1 Rejection-Sampled Supervised Fine-tuning
2.2.2 Rule-based Reinforcement Learning
2.3 Motivation
3 Approach
@@ 4.2 Metrics
To systematically evaluate the effectiveness of RepoSearcher with ToolTrain-model, we conducted a multi-dimensional evaluation using multiple metrics: Recall@k, MAP, MRR, nDCG@k and %Resolved. See Appendix E for detailed description of metrics.
@@ 7 Conclusion
In this paper, we improve the issue localization performance of LLMs by introducing ToolTrain, a tool-integrated training approach to enhance their tool-use and reasoning. Our framework first utilizes supervised fine-tuning to warm up the model with our lightweight agent, RepoSearcher, and then employs tool-integrated reinforcement learning to teach the model how to effectively navigate code repositories using various tools. We apply ToolTrain to open-source models, and evaluate them on issue localization tasks. The experiment results show that ToolTrain trained model achieve state-of-the-art performance among same-size models, and even surpassing leading models like Claude-3.7 on specific tasks. The code is available at https://github.com/Mizersy/RepoDeepSearch
@@ Appendix D Details of Baseline Frameworks
Appendix E Details of Metrics
Recall@k
MAP
[…]
MRR
[…]
nDCG@$k$
[…]
%Resolved